# WildAvatar: Web-scale In-the-wild Video Dataset for 3D Avatar Creation

### v2

@@ 3 WildAvatar
Our goal is to build a large-scale in-the-wild dataset for human avatar creation. To this end, we collect abundant real-world human videos from YouTube and annotate corresponding labels. Without time-consuming and high-cost manual filtering and annotation [15], we design an efficient pipeline to filter video candidates and obtain high-quality annotations. Specifically, we download human movement web videos with our automatic pipeline, as described in Sec. 3.1. Then, as described in Sec. 3.2, we perform several processing steps (e.g., filtering and annotation) on downloaded web videos. We collect a dataset of $10k+$ real-world human videos with $200k+$ frames and annotations.
3.1 Data Collection
We follow two steps to collect human movement videos: 1) We first search for a large diverse set of video candidates on YouTube. Note that video candidates could contain some noisy videos without human movement; 2) We then download candidate videos with automatic tools [3], cut them into clips, and filter out short video clips.
Downloading Video Candidates. To cover a wide range of in-the-wild human-central activities, we start from a label pool of human motion datasets [43]. Based on these human motion labels (e.g., fishing and playing tennis), we download over $100k$ video candidates from YouTube API.
Filtering Video Clips. Some collected video candidates could not meet the high-quality human avatar creation requirement. For example, human bodies may not exist (e.g., preview, occlusion) or frequently change poses (e.g., montage) in some frames of these videos. To filter out such videos, we cut video candidates into clips with SceneDetect [1] and discard video candidates with only short clips. After filtering video candidates, we obtain $460k$ video clip candidates for further processing.
3.2 Data Processing
To obtain high-quality annotations (i.e., SMPL, camera parameters, and human segmentation mask) for collected human videos, we adopt the following four stages to annotate human video clips.
Stage I: Human Bounding Box Detection and Tracking. Precise human bounding boxes are important for monocular SMPL estimation methods [36, 33, 41, 79] and human segmentation (using Segment Anything [34]). Hence, we first obtain the bounding box of human subjects with an off-the-shelf state-of-the-art Yolo detection method [55]. Then we filter out video clip candidates with either short (e.g., less than $2$ seconds) human activity sequences or low detection or tracking confidence. Finally, we only keep the tracked subject with the largest bounding box on average (i.e., key subject) over the whole video clip sequence.
Figure 2: The four-stage data processing pipeline. We first obtain the bounding box of key subjects in videos in Stage I and extract human segmentation masks in Stage II. Then the SMPL and camera parameters are coarsely estimated in Stage III and later refined in Stage IV.
Stage III: Coarse SMPL and Camera Estimation. We first estimate SMPL and camera parameters frame by frame, using a state-of-the-art single-image-based human pose and shape estimation method PARE [36]. Then, we check the pose and viewpoint shifts of the remaining video clips and only retain those with considerable viewpoint shifts and human movements. The estimated coarse SMPL and camera parameters will be used as initialization for implicit human reconstruction methods [72, 53, 51] to further refine SMPL and camera parameters in Stage IV.
Due to the depth ambiguity, monocular SMPL estimation and SAM might have low overlaps in estimated human masks. We double-check their consistency by comparing SMPL projection masks with the SAM masks. Intuitively, the SMPL mask denotes the naked body, while the SAM mask contains the clothed body. Therefore, the SMPL mask should be mostly covered by the SAM mask. To guarantee high-quality annotations, we discard video clips whose SAM-masked region has small overlaps with the SMPL-masked region.
Stage IV: Refining SMPL and Camera In-the-loop. The coarse SMPL estimated in Stage I may ignore the temporal consistency of human motions. Inspired by previous SMPL annotation pipelines [37, 2, 32], we further refine SMPL and camera parameters by smoothing these parameters across the whole video clip sequence via gradient descent. We also incorporate the estimated 2D keypoints and SAM masks into a refinement loop, providing additional supervision for precise SMPL and camera parameter estimation. Qualitative comparisons between coarse and refined SMPL and camera parameters can be found in the Fig. A1.
3.3 Data Analysis
@@
Our collected WildAvatar contains $10,647$ real-world human video clips with $200k+$ images. We randomly split WildAvatar into training, validation, and test splits with $7k$, $1.5k$, and $1.5k$ subjects, respectively. We show data statistics of our WildAvatar in Fig. 3 and describe the details of pose, viewpoint, and clothing distribution as follows.
@@
Poses: Various vs. Tedious. In Fig. 3 (c), we compare the pose distribution differences among WildAvatar and previous human datasets. Specifically, we visualize the top-2 components of body poses with t-SNE [54] dimension reduction. The pose distribution comparison shows that our WildAvatar contains more diverse poses than previous laboratory datasets, as laboratory datasets only contain several designed motion sequences. This illustrates the importance of our large-scale in-the-wild WildAvatar when applying the existing models to real-world scenarios.
@@
Viewpoints: Free vs. Fixed. We compare viewpoint distribution among previous human datasets [53, 9, 84] and WildAvatar. Specifically, we visualize the 3D rotations of observed cameras relative to human subjects. As suggested in Tab. 1, previous laboratory datasets are mainly collected in indoor lightstages, where the RGB(D) cameras’ positions are fixed for whole datasets. In real-world scenarios, however, the viewpoints of humans are arbitrary. As shown in Fig. 3 (d), the viewpoint distributions in previous datasets are sparse or unbalanced, while our WildAvatar covers a more dense viewpoint distribution.
@@
Cloth: Diversity vs. Unitary. Previous laboratory human datasets mainly contain unitary tight clothes, while complex real-world appearances (e.g., different hairstyles) and diverse loose clothes are seldom involved. Our WildAvatar involves diverse tight and loose clothes in the real world. To quantitatively measure the clothing diversity between real-world and laboratory data, we define a metric (SSIOU) that calculates the inverse IOU between SMPL [47, 50] projected masks and human segmentation masks among different datasets in Fig. 3 (f). For humans with tight clothes, the measured SSIOU is close to $1.0$ (blue in Fig. 3 (f)), indicating their SMPL masks are similar to human segmentation masks. Correspondingly, a larger SSIOU (red in Fig. 3 (f)) reveals that human clothes are looser, with segmentation masks being larger than SMPL masks. As shown in Fig. 3 (f), more than $30\%$ subjects in WildAvatar have SSIOU values over $1.4$, suggesting that WildAvatar covers various types of clothes. This analysis validates that WildAvatar satisfies data distributions in real-world scenarios.
@@ 4 Experiments
We evaluate previous state-of-the-art methods on our WildAvatar dataset with the following commonly used metrics (details in Sec. 4.1) on 3D avatar creation. To illustrate the unclaimed challenges for avatar creation in the wild, we benchmark previous popular per-subject methods in Sec. 4.2. To further demonstrate the benefit of the large-scale WildAvatar, we compare generalizable avatar creation methods in Sec. 4.3.
4.1 Evaluation Metrics
To quantitatively evaluate the quality of novel view and novel pose synthesis, we apply three commonly used metrics: peak signal-to-noise ratio (PSNR) [60], structural similarity index (SSIM) [71], and Learned Perceptual Image Patch Similarity (LPIPS) [80]. Following previous literature [38, 25], we compute whole-image metrics for per-subject reconstruction methods, while reporting the metrics based on projected 3D human bounding box areas for generalizable methods.
4.2 Track A: Per-Subject Human Avatar Creation
Table 2: Quantitative comparisons on WildAvatar benchmark. We report the quality of novel pose synthesis of popular methods, on both the lab benchmark ZJU-Mocap and our WildAvatar.
Method Train FPS WildAvatar ZJU-Mocap
@@ 4.2 Track A: Per-Subject Human Avatar Creation
NHR [73] 5h 1.23 18.72 91.1 10.1 28.25 89.2 12.2
NeuralBody [53] 10h 1.48 16.55 85.8 12.6 30.66 96.4 4.3
Animatable NeRF [51] 10h 1.11 19.43 91.0 10.4 29.77 96.5 4.7
Animatable SDF [52] 10h 0.40 19.20 91.9 11.4 30.38 97.5 3.7
HumanNeRF [72] 10h 0.30 22.65 87.1 15.4 30.66 96.9 3.3
InstantNVR [19] 5m 2.20 23.01 91.9 9.9 31.01 97.1 3.8
GauHuman [26] 1m 189 24.02 92.1 9.1 31.34 96.5 3.1
Figure 4: Qualitative comparisons of popular per-subject methods on WildAvatar. NHP and NB, AN, AS, HN, IN, and GH denote Neural Human Rendering, NeuralBody, Animatable NeRF, Animatable SDF, HumanNeRF, InstantNVR, and GauHuman, respectively. Faces are blurred to protect privacy.
Setup. To explore the quality of WildAvatar and investigate the maximum reconstruction performance on monocular in-the-wild videos. We evaluate popular per-subject avatar creation methods on our WildAvatar dataset. We evaluate a total of 7 baselines: Neural Human Rendering (NHR[73], NeuralBody (NB[53], Animatable NeRF (AN[51], Animatable SDF (AS[52], HumanNeRF (HN[72], InstantNVR (IN[19], and GauHuman (GH[26]. Considering the calculating and time cost, it is impractical and inefficient to train all scenes in WildAvatar. Instead, we manually select $10$ representative human subjects from WildAvatar for benchmarking. For each subject, we randomly choose $10$ frames for training and the remaining $10$ frames are for testing. We follow the default model settings (learning rate, batch size, etc.) in official implementations and report average metrics of these baselines in Tab. 2.
Quantitative and Qualitative Analysis. As shown in Tab. 2, existing methods struggle to handle the more challenging in-the-wild WildAvatar, resulting in unsatisfactory performance. Compared to the popular ZJU Mocap [53] benchmark, there is a marginal drop of PSNR ($-10.12$), SSIM ($-8.04$), and LPIPS ($+6.47$) on average. We further present qualitative comparisons in Fig. 4, and outline three potential challenges towards in-the-wild avatar creation in the following.
Noisy Annotation. The quality of avatar creation significantly depends on the quality of annotations. Although we streamline our annotation pipeline with the best existing methods, it is unrealistic to annotate in-the-wild scenarios as accurately as in the lab. Specifically, existing monocular SMPL estimation methods may fail in challenging areas (e.g., hands and feet), and segmentation methods may also convey unstable artifacts (See GT in Fig. 4 (f)). It is an imperative field that investigates methods for reconstructing the human body from noisy in-the-wild data.
Loose Cloth. In-the-wild data features a wide variety of clothing, rather than dominant tight garments (e.g., T-shirts and shorts) in existing laboratory datasets. These loose clothes introduce non-rigid motions and ambiguous deformation weights, which pose significant challenges for avatar creation methods. As shown in Fig. 4 (a), (e), and (f), existing methods struggle to reconstruct loose garments and are prone to render fuzzy or floating regions. These garments are positional ambiguous to render in novel poses, as they are not considered in the existing body models [47, 50].
Uneven Light Condition. Most state-of-the-art methods [51, 72, 26, 53] reconstruct avatars only in the canonical space, and they warp the space from canonical to observation in novel pose synthesis. This fashion assumes a uniformity of light distribution of the target scenes, which is easy to set up in laboratory lightstages. However, real-world scenarios are usually illuminated in partial-occluded directions. This uneven light condition makes it more challenging to learn diffuse-aware avatars from monocular videos. Current avatar reconstruction models struggle to distinguish the reflection effects and cloth textures. This ambiguity may lead to dissonant lighting effects, e.g., the dress region in Fig. 4 (d) and legs in Fig. 4 (f).
4.3 Track B: Generalizable Human Avatar Creation
Table 3: Generalizability comparisons on challenging WildAvatar and laboratory benchmarks. We report the results of previous generalizable avatar creation methods on the cross-domain setting. NP/NV denotes Novel Pose/View. RP/THU/HM/ZJU/WA are short for RenderPeople [56] / THuman [84] / HuMMan [9] / ZJU Mocap [53] / WildAvatar (Ours) dataset. Novel Lab Test refers to the test split of laboratory datasets not included in the training.
@@ 4.3 Track B: Generalizable Human Avatar Creation
Figure 5: Qualitative comparisons on the state-of-the-art generalizable avatar creation method [25]. “w/ WA” or “w/o WA” denotes training with or without WildAvatar, respectively. Faces are blurred to protect privacy.
@@
Setup. We evaluate four state-of-the-art generalizable baselines on our large-scale WildAvatar, including 1) Neural Human Performer (NHP[18], 2) MPS-NeRF [38], 3) SHERF [25], and 4) Generalizable 3D Gaussian Splatting (Generalizable GS). Specifically, the Generalizable GS is adapted from GauHuman [26], with the Gaussians’ attributes decoded from the input features, following the approaches in  [11, 45]. To evaluate the overall generalizability among various scenarios, we also report results on four laboratory datasets, including 1) ZJU-Mocap [53], 2) RenderPeople [56], 3) HuMMan [9], and 4) THuman [84]. We split these datasets into training and test sets as in SHERF [25]. To avoid overfitting to a small indoor dataset, we adopt the cross-domain testing strategy, e.g.we mask out the ZJU training set when testing on the ZJU.
@@
Quantitative and Qualitative Analysis. There are three main observations in Tab. 3. First, large-scale WildAvatar is essential for improving generalizability towards in-the-wild scenarios. Comparing odd and even rows, the WildAvatar training set brings considerable improvements in real-world scenarios. On the WildAvatar test set, it improves PSNR, SSIM, and LPIPS by 4.52%, 3.53%, and 5.06% on average, respectively. This suggests that using large-scale in-the-wild data is important for the generalizable avatar creation methods. As shown in Fig. 5, models without (w/o) training on WildAvatar, tend to predict dark-colored artifacts on the lower body. And models with (w/) training on WildAvatar can provide more realistic appearances. Second, large-scale in-the-wild data is also beneficial for laboratory benchmarks, although there are large domain gaps and the annotation of in-the-wild data is not as accurate as in the laboratory. By introducing WildAvatar dataset, we obtain an improvement of 2.57% and 2.50% PSNR on novel view and pose synthesis, respectively. Third, although 3D Gaussian Splatting (3DGS) performs well in per-subject human avatar creation, it shows slightly worse than NeRF in this generalizable setting. Our insight is that 3DGS highly relies on the point cloud initialization [26, 13, 45]. Unfortunately, due to the depth ambiguity, it is challenging to infer accurate point cloud initialization from a single image.
5 Conclusions
@@
We curate WildAvatar, a large-scale in-the-wild human avatar creation dataset from YouTube, with 10, 000+ various human subjects and scenes. Compared with traditional avatar creation datasets, our WildAvatar consists of at least 10×10$\times$10 × more subjects or scenes. We benchmark state-of-the-art human avatar creation methods on WildAvatar, illuminate the potential of the avatar creation task with data at scale, and reveal its remaining challenges towards real-world applications. We hope this dataset will shed insights on following in-the-wild human avatar creation methods, and benefit 3D/4D human content generation works.
Pipeline Efficiency. The efficiency of data collection and annotation is also crucial to facilitate data scale-up and application. Despite the complex multiple-stage design, our data-collecting pipeline only takes less than $200$ seconds to generate full annotations for an arbitrary in-the-wild video.
Limitations. Despite the advantages of WildAvatar, we would like to list several limitations: 1) WildAvatar focuses on the 3D modeling of human bodies, lacking the 3D information of background environments. 2) To meet the needs of the avatar creation, WildAvatar contains video scenes involving dynamic human motions. Human poses in static activities (e.g., reading, fishing) are rarely included. 3) Multi-human interactions within the identical scenes are not covered.
Societal Impacts. In-the-wild avatar creations could enable potentially harmful uses. For instance, the unauthorized creation of avatars may violate portrait rights. Additionally, models trained on WildAvatar might exhibit biases in YouTube videos. Users should consider this when deploying, analyzing, and developing these models.
Appendix A Details in Data Pipeline

### v4

@@ 4 WildAvatar
Our collected WildAvatar contains $10,647$ real-world human video clips. We randomly split WildAvatar into training, validation, and test splits with $7k$, $1.5k$, and $1.5k$ subjects, respectively. Notably, although selected by our protocols, the remaining curated data is still much more diverse than in the laboratory systems (See Section 4 for statistics and Fig. A for unusual camera viewpoints and body poses). We show data statistics of our WildAvatar in Fig. 4 and describe the details of pose, viewpoint, and clothing distribution as follows.
@@
Poses: Various vs. Trivial. In Fig. 4 (c), we compare the pose distribution differences among WildAvatar and previous human datasets. Specifically, we visualize the top-2 components of body poses with t-SNE [62] dimension reduction. The pose distribution comparison shows that our WildAvatar contains more diverse poses than previous laboratory datasets, as laboratory datasets only contain several designed motion sequences. This statistic illustrates the importance of our large-scale in-the-wild WildAvatar when applying the existing models to real-world scenarios.
@@
Viewpoints: Free vs. Fixed. We compare viewpoint distribution among previous human datasets [60, 9, 100] and WildAvatar. Specifically, we visualize the 3D rotations of observed cameras relative to human subjects. As suggested in Table 1, previous laboratory datasets are mainly collected in indoor lightstages, where the RGB(D) cameras’ positions are fixed for whole datasets. In real-world scenarios, however, the viewpoints of humans are arbitrary. As shown in Fig. 4 (d), the viewpoint distributions in previous datasets are sparse or unbalanced, while our WildAvatar covers a more dense viewpoint distribution.
@@
Cloth: Diversity vs. Unitary. Previous laboratory human datasets mainly contain unitary tight clothes, while complex real-world appearances (e.g., different hairstyles) and diverse loose clothes are seldom involved. Our WildAvatar involves diverse tight and loose clothes in the real world. To quantitatively measure the clothing diversity between real-world and laboratory data, we define a metric (SSIOU) that calculates the inverse IOU between SMPL [55, 58] projected masks and human segmentation masks among different datasets in Fig. 4 (f). For humans with tight clothes, the measured SSIOU is close to $1.0$ (blue in Fig. 4 (f)), indicating their SMPL masks are similar to human segmentation masks. Correspondingly, a larger SSIOU (red in Fig. 4 (f)) reveals that human clothes are looser, with segmentation masks being larger than SMPL masks. Examples of various SSIOU can be found in Fig. D. As shown in Fig. 4 (f), more than $30\%$ subjects in WildAvatar have SSIOU values over $1.4$, suggesting that WildAvatar covers various types of clothes. This analysis validates that WildAvatar more satisfies data distributions in real-world scenarios.
@@ 5 Experiments
Method HMR2.0 Ours
@@
NHR [85] 18.23 89.7 10.3 18.93 91.6 9.9
NB [60] 16.17 84.2 12.6 16.74 86.2 12.1
AN [59] 19.02 89.1 11.4 19.72 90.5 10.9
AS [61] 18.72 90.9 11.1 19.20 92.3 10.7
HN [84] 22.52 86.3 15.2 23.12 88.0 14.6
IN [22] 24.39 92.7 8.2 25.28 94.3 7.7
GH [29] 24.73 93.8 6.3 25.89 95.7 5.7
Table 4: Quantitative comparisons on WildAvatar. We report the quality of novel pose synthesis of popular methods on the state-of-the-art HMR2.0 annotations and our annotations, revealing the advantages of our pipeline towards in-the-wild avatar creation.
Figure 5: Qualitative comparisons of avatars created with our annotations and the state-of-the-art HMR2.0. IN and GH denote InstantNVR and GauHuman, respectively. More accurate avatars can be created with our annotations. Human faces are blurred.
We conduct exploratory experiments on our WildAvatar with the following commonly used metrics (details in Section 5.1) on 3D avatar creation.  Section 5.2 illustrates the real-world avatars from our WildAvatar, created by per-subject avatar creation methods. It also illustrates the improvement of our annotating pipeline, by comparing annotations from our pipeline or the state-of-the-art HMR2.0 [23].  Section 5.3 further demonstrates the potential of generalizable avatar creation methods, when provided our large-scale WildAvatar.
@@
Table 5: Generalizability comparisons on challenging WildAvatar and laboratory benchmarks. We report the results of previous generalizable avatar creation methods on the cross-domain setting, offering a clear perspective across different domains. NP/NV denotes Novel Pose/View. RP/THU/HM/ZJU/WA are short for RenderPeople [64] / THuman [100] / HuMMan [9] / ZJU Mocap [60] / WildAvatar (Ours) dataset. Novel Lab Test refers to the test split of laboratory datasets not included in the training.
5.1 Evaluation Metrics
To quantitatively evaluate the quality of created avatars, we apply three commonly used metrics: peak signal-to-noise ratio (PSNR) [68], structural similarity index (SSIM) [83], and Learned Perceptual Image Patch Similarity (LPIPS) [97]. Following previous literature [45, 31], we compute whole-image metrics for per-subject reconstruction methods, while reporting the metrics based on projected 3D human bounding box areas for generalizable methods.
5.2 Per-Subject Avatar Creation
Setup. To explore the quality of WildAvatar and investigate the reconstruction performance on monocular in-the-wild videos with our annotations. We evaluate popular per-subject avatar creation methods on our WildAvatar dataset. We consider a total of 7 baselines: Neural Human Rendering (NHR[85], NeuralBody (NB[60], Animatable NeRF (AN[59], Animatable SDF (AS[61], HumanNeRF (HN[84], InstantNVR (IN[22], and GauHuman (GH[29]. Considering the calculating and time cost, it is impractical and inefficient to train all subjects and scenes in WildAvatar. Instead, we manually select $133$ representative human subjects from WildAvatar for exploring. For each subject, we randomly choose $100$ frames for training and $100$ for testing. We follow the default model settings (learning rate, batch size, etc.) in their official implementations and report average metrics of these baselines in Table 4.
Quantitative and Qualitative Analysis. As shown in Table 4, more accurate avatars can be created with our annotations. Compared to the annotations from the state-of-the-art HMR2.0 [23], our annotations improve PSNR ($+3.51\%$), SSIM ($+1.91\%$), and LPIPS ($-5.06\%$) on average. We further present qualitative comparisons in Fig. 5. Avatars from our annotations have more accurate extremities (e.g., feet, arms, and heads), which are challenging to predict and align from single images. This improvement may prove the benefit of Stage IV (e.g., 2D alignment and temporal smoothness).
5.3 Generalizable Avatar Creation
@@
Figure 6: Qualitative comparisons on the state-of-the-art generalizable avatar creation method [31]. “w/ WA” or “w/o WA” denotes training with or without WildAvatar, respectively. Human faces are blurred to protect privacy.
@@
Setup. We evaluate four state-of-the-art generalizable baselines on our WildAvatar, including 1) Neural Human Performer (NHP[21], 2) MPS-NeRF [45], 3) SHERF [31], and 4) Generalizable 3D Gaussian Splatting (Generalizable GS). Specifically, the Generalizable GS is adapted from GauHuman [29], with the Gaussians’ attributes decoded from the input features, following the approaches in  [12, 53]. To evaluate the overall generalizability among various scenarios, we also report results on four laboratory datasets, including 1) ZJU-Mocap [60], 2) RenderPeople [64], 3) HuMMan [9], and 4) THuman [100]. To avoid overfitting to a small indoor dataset, we adopt the cross-domain testing strategy, e.g., mask out the ZJU training set when testing on the ZJU.
@@
Quantitative and Qualitative Analysis. There are three main observations in Tab. 5. First, large-scale WildAvatar is essential for improving generalizability towards in-the-wild scenarios. Comparing odd and even rows, the WildAvatar training set brings considerable improvements in real-world scenarios. On the WildAvatar test set, it improves PSNR, SSIM, and LPIPS by 4.52%, 3.53%, and 5.06% on average, respectively. Notably, this is fairly evident as over $60\%$ of pixels are background. This improvement suggests that large-scale in-the-wild data is essential for the generalizable avatar creation methods. As shown in Fig. 6, models without (w/o) training on WildAvatar, tend to predict dark-colored artifacts on the lower body. Models with (w/) training on WildAvatar can provide more realistic appearances and geometries. Second, large-scale in-the-wild data is also beneficial for laboratory benchmarks. However, there are significant domain gaps, and the annotation of in-the-wild data is not as accurate as in the laboratory. By introducing WildAvatar dataset, we obtain an improvement of 4.17% and 4.54% PSNR on novel view and pose synthesis, respectively. Third, although 3D Gaussian Splatting (3DGS) performs well in per-subject human avatar creation, it shows slightly worse than NeRF in this generalizable setting. Our insight is that 3DGS highly relies on the point cloud initialization [29, 15, 53]. Unfortunately, due to the depth ambiguity, it is challenging to infer accurate point cloud initialization from a single image.
6 Conclusion
@@
This work introduces an automatic pipeline with filtering protocols for collecting in-the-wild human annotations from the web. From this pipeline, we curate WildAvatar, a large-scale in-the-wild human avatar creation dataset from YouTube, with $10,000+$ various human subjects and scenes. Compared with traditional avatar creation datasets, our WildAvatar consists of at least 10×10$\times$10 × more subjects or scenes. We demonstrate the effectiveness of the proposed pipeline and illuminate the potential of WildAvatar on avatar creation tasks with data. We hope our pipeline and dataset will shed insights on following in-the-wild human avatar creation and benefit 3D/4D human content generation works.
Acknowledgement. This study is supported under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).
Appendix A Details in Data Collection
Downloading Video Candidates. Our first goal is to collect videos from the web with human motions. To cover a wide range of in-the-wild human-central activities, we start from a label pool of human motion datasets [51]. Based on these human motion labels (e.g., fishing and playing tennis), we download over $100k+$ video candidates from YouTube API.
Per-filtering Video Clips. Some collected video candidates could not meet the high-quality human avatar creation requirement. For example, human bodies may not exist at all (e.g., blank preview, severe occlusion) or frequently change across scenes (e.g., montage) in some subsections of these videos. To exclude such unqualified subsections, we utilize SceneDetect [1] to cut video candidates into clips and eliminate those with insufficient length (less than $2$ seconds). Subsequently, we apply human detection models with low FPS to these clips to efficiently filter out those without human subjects at minimal cost. After filtering video candidates, we obtain $460k+$ video clip candidates for further processing.
Appendix B Details in Data Pipeline